OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, -3) and (3, -5). 1 1/2 0 -1

OpenStudy (anonymous):

Please Help! My guess is 1/2 but I'm not sure... :/ I want to make sure I have the right answer before I continue...

OpenStudy (onepieceftw):

\[y_1-y_2/x_1-x_2\]

OpenStudy (onepieceftw):

That to calculate slope

OpenStudy (onepieceftw):

So -3-(-5)/1-3 What does that equal?

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (1,-3)...x1 = 1 and y1 = -3 (3,-5)....x2 = 3 and y2 = -5 now we sub slope = (-5 - (-3) / (3 - 1) slope = (-5 + 3) / 2 can you finish ?

OpenStudy (anonymous):

The answer is -1, right? That's what I got @OnePieceFTW @texaschic101

OpenStudy (onepieceftw):

That's right.